OpenStudy (anonymous):

Please help me (2sqrt5 + 3sqrt7)^2

OpenStudy (anonymous):

\[(2\sqrt{5} + 3\sqrt{7})^{2}\]

OpenStudy (anonymous):

Use the formula : \[( a + b )^{2} = a ^{2} ++2ab + b ^{2} \]

OpenStudy (anonymous):

And then plug in the values: \[a = 2 \sqrt{5}\] \[b = 3\sqrt{7}\]

OpenStudy (anonymous):

Now, you get : \[(2\sqrt{5})^{2} + 2 \times 2\sqrt{5}\times 3\sqrt{7} + (3\sqrt{7}) ^{2}\]

OpenStudy (anonymous):

Can you do it now?

OpenStudy (anonymous):

I have a real hard time simplifying further than that... I can't really do radicals

OpenStudy (anonymous):

Okay, well take \[(2\sqrt{5} ) ^{2}\] separately then. It'll be equal to \[2 \times \sqrt{5} \times \sqrt{5}\] How much is it?

OpenStudy (anonymous):

Make the 2 into 4. Oops. :P

OpenStudy (radar):

Do it the old fashion way

OpenStudy (radar):